AGW3 Alignment for LUCI1 - 20180603 UT

Observers: JHill, BRothberg (Tucson)
Telescope Operator: SAllanson (LBTO)
AO Support: GTaylor (Tucson)
Telescope Support: MikeW (LBTO)
Software Support:
Instrument Support: ESolheid (LBTO)
Engineering Support: DCox


Functional Checks on the Telescope (completed 20180602 UT)

- DONE probe motion with setxy/getxy

- DONE wfs and guide camera images

- DONE telescope authorization with GCS

- DONE test preset (w/o fake guide stars)

- DONE check WFS_lenslet_edge


Sunset - 02:33
Twilight End - 04:07
Night Center - 07:25
Twlight Start - 10:42
Sunrise - 12:16

Moon ~ 81%

Plan for On-sky Checks and Calibrations

DONE AO Checkout in Twilight (DX only)

DONE Verify that guiding works as expected with IE or CA perturbations. [0.2 hour]

Adjust guidecam_hotspot_x,_y (BGTechProcedures) [0.3 hour] - We never managed to detect light on WFS, See IT 7282

Measure on-axis pupil position and update GCS and IDL parameters. [0.5 hour] WFScam_hotspot_x,_y Use gcs_wfsc_display and GCSGUI to check.

Iterate guidecam_hotspot and WFScam_hotspot with collimation.

Set Probe center position in AGW3 (BGTechProcedures)[0.5 hour, prefer good seeing]

Skip Pupil Wobble measurement as SX M3 mounting is unchanged, but beware IT 7269. (Commissioning/ComTools#PupilWobble)

Check off-axis guiding and collimation. [0.2 hour]

Check LUCI dithers. [0.2 hour] - longslit and imaging

Check LEFTZEROPOINT wrt/LUCI slit [0.5 hour] - MOS masks StoneJ or StoneL

Check location of rotator center on LUCI1. [0.5 hour]

Collect LUCI1 and LUCI2 transform data [2 hours, requires good seeing] (We are not anticipating data reduction during the night.)

In case of bad seeing (> 1.5"):
  • (D-time) H-band spectral library, seeing >>2" - many mags extinction - LUCI-Bino G210-H (Barry)
  • (D-time) Near-IR Local Merger spectroscopy zJHK G200 - LUCI-Bino
  • (E-time) TMS Active with LBC (John H. and Andrew) (seeing <2", up to 3 mags clouds, target flexible)
  • (E-time) LUCI+MODS Binocular preset tests ( IT 6343) - has not been addressed in over a year, but the IT suggests the fix needs to be tested (John H. and Barry) Target: IC 5117 ~ 10min
  • (E-time) LBC+MODS pseudo-monocular test ( IT 5715) - verify that LBC will dither without streaking in exposures (John H. and Barry) Target: Hoag's Object ~ 10min
  • (E-time) LBC - WRS testing on tech chips
  • (E-time) MODS AGW transforms (blue channel both sides)
Any Conditions:
  • (E-time) obtain zeropoint measurements at J-band for LUCI-1 and LUCI-2 (seeing limited) for Arcetri (goal to constrain where the flux loss is coming from in the optical path).
D-time fill in programs:
  • DONE Dwarf Galaxy Mergers - LUCI-Bino (LUCI-2 in case of AGW-3 issues)


  • AGW3 and LUCI1 mounted on 20180601 daytime.
  • GCS software configuration updated as per IT 7252.
  • AGW3 functional checkout completed.


AO Checkout

UT 01:09 - Starting up LUCIs and Initializing all

UT 01:10 - Test darks taken LUCI-1 and LUCI-2 luci1.20180603.0001-0002 luci2.20180603.0001-0002

UT 01:29 - Configuring LUCI-2 for AO setup (N30/fieldstop)

UT 01:46 - luci2.20180603.0003-0009 - test setup AO checkup

UT 02:13 - Enclosure opening

UT 02:39 - moving to a pointing star, then collimation star - will begin AO on-sky checkout

UT 02:56 - progressing through checkout

UT 03:01 - Greg sent FeII and clear but received an instrument gui error and it went to OH1060

UT 03:03 - error on filter wheel 1 (narrow filters), running init via engineering gui panel - fixed it

UT 03:04 - Running Doug's script for taking data - luci2.20180603.0010-0012 - AO LUCI images

UT 03:07 - re-authorizing binocular LUCI - shutting down wfs, putting in blinds in LUCI-2

AGW3 alignment in front of LUCI1

Doug (in London) has measured lenslet_x_center to be 56.0,145,5 (consistent with previous AGW3 values). I've added 13.3*8.5 to get updated GCS values for WFS_lenslet_edge_x of 169.0,258.5

UT 03:15 AGW3 MOXA communication is hung. Steve is going to cycle power. Then rebooting oac.

UT 03:30 relax maxOffsetToHotspot from 0.54 to 2.54 arcsec in preparation for guiding hotspot alignment.

UT 03:45 Had to do this to select AGW: gcsclient selectAGW left LUCI_L.cfg

UT 03:50 Repeat selectAGW after putting WFScam_hotspot_sext_x _y back in LUCI_L.cfg (They cannot be commented out.)

UT 03:54 ACQUIRE preset. Left Acquire image is flipped in X=CA for Parallactic mode preset. Is there a GCSGUI config file that knows about AGW3? SX IE=-25 CA=+42 DX IE=-33 CA=-36

UT 04:01 Home SX M3 while watching Acquire movie to check for IT 7269. Star came right back to the same spot.

04:08 ACTIVE prset -- Acq offset went the wrong way in X.

UT 04:11 pointingCorr _flip_x to -1 guidingCorr_flip_x to -1

04:13 manually rastering hotspot, but can't find it. Nothing but noise on wfs image.

04:29 ACTIVE preset to 3rd mag star too bright


05:00 100sec wfs exposure with Jupiter on acquire - no photons detected by wfs

05:03 Steve cycling AGW3 power and restarting GCSL.

wfsc000006 still nothing on wfs

Abandon ship and switch over to Barry's dwarf galaxy program.



First target is SDSSJ 1416-0127 w/LUCI-2

luci2.20180603.0013-0016 acq, 0017-0040 spectra

UT 05:20 - configuring for LUCI-2 only

UT 05:21 - sending preset

UT 05:37 - starting science DIMM 0.8"

UT 05:45 - can see both galaxies in the slit, nice bright Pa alpha, Br Gamma, continuum is a bit fainter than expected...but super cool!

UT 06:14 - completed 1800sec, running a 2nd round since we are in pirate mode

UT 06:52 - completed SDSSJ 1416-0127 - moving to Telluric HD 123233

UT 06:56 - HD 123233 luci2.20180603.0041-0045

UT 06:59 - starting telluric integrations

UT 07:02 - Done


UT 07:03 - moving to telluric for SDSSJ 1518+1440 (HD 124773) luci2.20180603.0046-0051

UT 07:10 - slit centering slightly off, offset +0.36" DETXY

UT 07:11 - starting integration

UT 07:15 - sending preset to SDSSJ 1518+1440 luci2.20180603.0052-0085 (1hr 15min)

UT 07:18 - acqusition went to H-filter, not sure why, scripts had Ks for acquisition images like previous target

UT 07:22 - RTD is super slow for pair subtraction again

UT 07:33 - starting integration - hard to tell - we'll see in 5 minutes

UT 07:43 - def see one bright Pa alpha line (some extension?) first exposure (0056 has persistence from slit - might need to ditch it)

UT 08:13 - finished first run through script, running it again (since we're in pirate mode)

UT 08:55 - adding 15 more minutes (6 more exposures)

UT 09:10 - finsihed target


UT 09:11 - moving to SDSSJ 1614+4101 - luci2.20180603.0086-0119

UT 09:13 - missed the guidestar - just off to the right - Steve nudging the star over

UT 09:13 - sending preset again

UT 09:27 - confirmed via IRAF/DS9 that blob is in there - starting spectroscopy

UT 09:34 - persistence from slit again - likely need to ditch the first exposure

UT 09:46 - DIMM ~1.1" Guider ~1.0"

UT 09:44 - seeing bubbling up a bit - good thing the REALLY faint one was done in <1" seeing

UT 09:49 - seeing bubble 1.5" -0097

UT 10:01 - DIMM up to 1.6" but guider showing ~1"

UT 10:04 - finished first set, running script again

UT 10:42 - comlpeted set, doing 6 more exposures (total of 1hr 15min on-target)

UT 11:00 - completed SDSSJ 1614+4101

UT 11:01 - moving to Telluric HIP 87643 luci2.20180603.0120-0124

UT 11:08 - starting telluric spectral exposures

UT 11:12 - Closing up - finished telluric - preparing for calibratiosn


LUCI-2 G200 HKspec 1pt5" slit cals

UT 11:25 -starting cals

0125-0129 - flat lights off

0130-0134 - flat lights on

0135-0136 - arc lights off

0137-0138 - argon lights on

0139-0140 - xenon lights on

0141-0142 - neon lights on


LUCI-1 zJspec G200 0.5" slit / LUCI-2 HKspec G200 0.5" slit

-had to remake OT script to allow mixed grating configurations on each side

L1#0014-0018 flat lights off, 0019-0021- BAD AGW on-axis, need to redo

L1#0023-0027 flat-lights off 0028-0032 flat lights on, 0033-0034 Argon, 0035-0036 Neon, 0037-0038 60sec arc lights off

L2#0143-0147 flat lights off, 0148-0152 flat lights on, 0153-0154 arc lights off, 0155-0156 Argon, 0157-0158 Xenon, 0159-0160 Neon


LUCI-1 zJspec G200 0.75" slit / LUCI-2 HKspec G200 0.75" slit

L1#0039-0043 flat lights off, 0044-0048 flat lights on, 0049-0050 arc lights off, 0051-0052 Argon, 0053-0054 Neon

L2#0161-0165 flat lighst off, 0166-0170 flat lights on, 0171-0172 arc lights off, 0173-0174 Argon, 0175-0176 Xenon, 0177-0178 Neon

UT 12:30 - putting LUCI-2 into safe mode (LUCI-1 still going)

UT 12:43 - putting LUCI-1 into safe mode

Done for this morning


Daytime testing of AGW3 WFS

UT 19:24 Authorize as LUCI/LUCI, setxy to move AGw3 and AGW2 guide probes to center of the field. All dome lights are on as LBTI folks are working on SOUL installation.

UT 19:26 gcsclient readWFSCam left -e 1000 ---> /lbt/data/share/tcs/lbto.20180603/left_wfsc000008.fits (1 sec exposure has only uniform bias level at 6150 ADU, bias has a normal looking bias structure, but no indication of an image)

UT 19:27 gcsclient readWFSCam right -e 1000 ---> /lbt/data/share/tcs/lbto.20180603/right_wfsc000001.fits (1 sec nice image with bias of 580 and 1000 ADU in the lenslet circles)

UT 19:38 gcsclient readWFSCam left -e 100000 ---> /lbt/data/share/tcs/lbto.20180603/left_wfsc000009.fits (100 sec exposure has only uniform bias level at 6150 ADU, bias has a normal looking bias structure, but no indication of an image)

TIME-OBS= '19:39:40.976'       / Date of observation - UTC GCS                  
DATE-OBS= '2018-06-03'         / Universal Time - UTC GCS                       
AGWTYPE = 'AIP_LUCIFER'        / type of AGw unit used to acquire image         
AGWNAME = 'AIPAGW3'            / unique name to identify the AGw unit           
CCDTEMP =           -999.90002 / last returned CCD temperature                  
DEWRTEMP=           -999.90002 / last returned dewar temperature                

UT 19:46 - Authorize as MODS/MODS

UT 19:47 stopAGW -u 3 (power off)

UT 19:51 startAGW -u 3 (power on)

UT 19:53 - Authorize as LUCI/LUCI (GCS connected to both AGw correctly)

UT 19:57 setxy -u 3 -x 0 -y 612.0

UT 19:58 gcsclient readWFSCam left -e 1000 ---> /lbt/data/share/tcs/lbto.20180603/left_wfsc000010.fits

UT 20:00 gcsclient readWFSCam left -e 10 ---> /lbt/data/share/tcs/lbto.20180603/left_wfsc000011.fits

UT 20:01 gcsclient readWFSCam left -e 100000 ---> /lbt/data/share/tcs/lbto.20180603/left_wfsc000012.fits

-- %USERSIG{JohnHill - 2018-05-31}%
Topic revision: r13 - 21 Jun 2018, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback